OpenStudy (anonymous):

Solving Linear Systems With Three Variables. Solve. Please show steps. {x+5y-2z=15 {3x-7y+4z=-7 {2x+4y-3z=21

OpenStudy (unklerhaukus):

\[\left[\begin{matrix}1&5 & -2 \\ 3 &-7 &4\\2 &4&-3 \end{matrix}\right]\left(\begin{matrix} x\\y \\ z\end{matrix}\right)=\left(\begin{matrix} 15\\-7 \\21 \end{matrix}\right)\]

OpenStudy (radar):

2x+10y-4z=30 2x+4y-3z= 21 -------------- 6y -z =9 3x+15y-6z=45 3x-7y +4z=-7 -------------- 22y-10z=52 60y-10z=90 -------------- -38y =-38 y=1 6(1) -z=9 6-z=9 -z=9-6=3 z=-3 2x+4(1)-3(-3)=21 2x+4+9=21 2x=21-4-9=8 x=4 y=1 z=-3